A family is driving from Vancouver to Kelowna, a distance of 389.5 km. During the first 200 km, they average 103 km/h. They stop for lunch for 67 min. They continue driving and get to Kelowna 3 h and 47 min later. What was their average speed (in km/h) for the entire journey, including the break?
1
Expert's answer
2020-11-23T10:27:16-0500

Their average speed for the entire journey:


"V=\frac{S}{T}\\V=\frac{389.5}{\frac{200}{103}+\frac{67}{60}+\frac{3(60)+47}{60}}\\V=57\frac{km}{h}"